# B-cos LM: Efficiently Transforming Pre-trained Language Models for Improved Explainability

### v1

@@ Appendix F SeqPG Example
Figure 7 presents a SeqPG example from AG News using B-cos BERT. For better visualization, each segment is truncated to 20 tokens instead of 50 used in the experiments. Unlike the hybrid document evaluation proposed by Poerner et al. (2018), our approach explicitly controls segment length and position to ensure a fair comparison. Additionally, we measure the proportion of correctly assigned positive attributions rather than relying solely on the highest attribution value.
Appendix G Task Performance of Other B-cos LMs
Figure 8: Mean accuracy of conventionally fine-tuned and B-cos DistilBERT models averaged over three runs. B-cos models perform comparably to conventional models on most tasks.
Figure 9: Mean accuracy of conventionally fine-tuned and B-cos RoBERTa models averaged over three runs. B-cos models perform comparably to conventional models on most tasks.
@@ Appendix G Task Performance of Other B-cos LMs
Appendix H Faithfulness Evaluation of Other B-cos LMs
Tables 6 and 7 present the faithfulness evaluation results for DistilBERT and RoBERTa. The findings are consistent with our main experiments (cf. Table 2), confirming that B-cos LMs produce more faithful explanations compared to post-hoc explanation methods.
@@ Appendix H Faithfulness Evaluation of Other B-cos LMs
(a) Conv. DistilBERT Attention 26.36 5.37 50 31.62 10.46 50 30.56 14.67 50
@@
(b) B-cos DistilBERT Attention 28.47 3.05 50 31.36 4.15 50 37.33 6.49 50
IxG 22.33 9.09 58.63 51.02 -1.44 53.76 41.62 0.29 56.03
SIG 14.73 5.62 53.09 39.75 -0.11 64.18 28.68 7.27 60.75
DecompX - - - - - - - - -
ShapSampl 31.78 1.77 62.60 64.65 -2.42 56.89 34.64 4.56 55.8
LIME 58.25 0.31 77.65 69.96 -0.43 61.08 44.66 1.66 59.27
(c) B-cos DistilBERT B-cos 61.93 -1.01 86.78 75.73 -2.57 71.95 57.2 -4.49 74.89
Table 6: Faithfulness evaluation for conventionally fine-tuned DistilBERT and B-cos DistilBERT across three datasets. The best results are in bold. We find that B-cos explanations are consistently more faithful than post-hoc explanations from both models.
@@
(a) Conv. RoBERTa Attention 22.17 3.80 50 25.26 5.84 50 32.94 7.52 50
@@
(b) B-cos RoBERTa Attention 16.07 6.83 50 29.83 2.85 50 27.35 8.39 50
IxG 22.25 2.39 56.15 67.2 -2.26 56.95 40.69 -1.11 58.59
SIG 44.35 -0.95 51.70 74.70 -2.39 58.03 51.20 -5.80 57.62
DecompX - - - - - - - - -
ShapSampl 55.26 -0.92 72.65 74.3 -2.39 62.74 51.54 -5.64 70.58
LIME 23.01 2.46 63.11 37.14 0.78 53.15 29.86 5.74 63.61
(c) B-cos RoBERTa B-cos 62.47 -1.18 86.63 75.15 -2.39 75.83 51.33 -5.18 74.01
@@
Table 7: Faithfulness evaluation for conventionally fine-tuned RoBERTa and B-cos RoBERTa across three datasets. The best results are in bold. We find that B-cos explanations are consistently more faithful than post-hoc explanations from both models.
Appendix I Human Evaluation Details
@@
In the human study, we select only examples shorter than 25 tokens for HateXplain and 40 tokens for AG News to improve visualization. Additionally, we replace [CLS] and [SEP] with ## to make the examples more understandable for lay users. Below, we provide the instructions along with a detailed description of the criteria and scoring used in our human evaluation.
@@
WARNING: SOME CONTENT IN THIS QUESTIONNAIRE IS HIGHLY OFFENSIVE. Prerequisites: Proficiency in English is required for this evaluation task. If you do not meet this criterion, please do not proceed. We invite you to review 50 examples where NLP models perform classification tasks and provide explanations for their predictions. The first 25 examples come from a hate speech detection task, where the model predicts whether a text is toxic or not toxic. The last 25 examples come from a topic classification task, where the model categorizes a text into one of four topics: sports, world, business, or sci/tech. For each example: The model’s prediction is shown along with four explanations justifying the prediction. The order of the explanations is randomized to prevent bias. Words highlighted in green indicate words that had a positive influence on the prediction, while words in red indicate words that had a negative influence. The intensity of the color reflects the strength of the impact. Important: The model’s prediction may be incorrect. Your task is to evaluate the explanations based on how well they support the model’s prediction, not the true labels. Evaluation Task: After reviewing each example, please rate the the human interpretability and human agreement of the four explanations on a scale of 1 to 5. Refer to the definitions and rating scales provided below when making your assessments.
@@
Appendix J More Examples of B-cos Explanations
We provide two more examples of B-cos and other (post-hoc) explanations from AG News in Figure 12. Consistent with our findings in § 6, B-cos LMs provide more human interpretable explanations.
@@
Appendix K Impact of B on Input-weight Alignment
@@
Appendix L Effects of B on Other Metrics
@@
Appendix M B-cos Explanations with Different B Values
@@
Appendix N Example of Model Bias

### v2

@@ Appendix D SeqPG Example
Figure 7 presents a SeqPG example from AG News using B-cos BERT. For better visualization, each segment is truncated to 20 tokens here instead of 50 used in the experiments. Unlike the hybrid document evaluation proposed by Poerner et al. (2018), our approach explicitly controls segment length and position to ensure a fair comparison. Additionally, we measure the proportion of correctly assigned positive attributions rather than relying solely on the highest attribution value.
Appendix E Task Performance of Other B-cos LMs
Figure 8: Mean accuracy of conventionally fine-tuned and B-cos DistilBERT models averaged over three runs. We use B=1.5, 1.25, and 1.5 for AG News, IMDB and HateXplain, respectively. B-cos models perform comparably to conventional models on most tasks.
Figure 9: Mean accuracy of conventionally fine-tuned and B-cos RoBERTa models averaged over three runs. We use B=1.5, 1.25, and 1.5 for AG News, IMDB and HateXplain, respectively. B-cos models perform comparably to conventional models on most tasks.
@@ Appendix E Task Performance of Other B-cos LMs
Appendix F Faithfulness Evaluation of Other B-cos LMs
Tables 9 and 10 present the faithfulness evaluation results for DistilBERT and RoBERTa. The findings are consistent with our main experiments (cf. Table 2), confirming that B-cos LMs produce more faithful explanations compared to post-hoc explanation methods.
@@ Appendix F Faithfulness Evaluation of Other B-cos LMs
Conv. DistilBERT Attention 26.36 5.37 50 31.62 10.46 50 30.56 14.67 50
@@
B-cos DistilBERT B-cos 61.93 -1.01 86.78 76.26 -1.28 72.68 57.2 -4.49 74.89
Table 9: Faithfulness evaluation for conventionally fine-tuned DistilBERT and B-cos DistilBERT across three datasets. We use B=1.5, 1.25, and 1.5 for AG News, IMDB and HateXplain, respectively. The best results are in bold. We find that B-cos explanations are consistently more faithful than post-hoc explanations from both models.
@@
Conv. RoBERTa Attention 22.17 3.80 50 25.26 5.84 50 32.94 7.52 50
@@
Saloss RoBERTa Attention 40.69 2.77 50 24.51 4.33 50 47.04 7.83 50
IxG 6.19 27.30 52.46 10.98 12.30 47.92 22.78 25.78 49.49
SIG 6.91 27.22 56.84 11.53 13.76 62.10 43.77 5.02 58.67
DecompX 61.46 0.16 74.20 65.50 0.10 74.41 54.94 2.47 65.63
ShapSampl 34.48 0.73 64.67 48.53 0.82 63.04 55.80 1.49 64.53
LIME 15.93 8.03 55.17 18.04 6.47 50.94 29.62 15.78 56.00
B-cos RoBERTa B-cos 62.47 -1.18 86.63 73.87 -2.30 74.05 51.33 -5.18 74.01
@@
Table 10: Faithfulness evaluation for conventionally fine-tuned RoBERTa, Saloss RoBERTa and B-cos RoBERTa across three datasets. We use B=1.5, 1.25, and 1.5 for AG News, IMDB and HateXplain, respectively. The best results are in bold. We find that B-cos explanations are consistently more faithful than post-hoc explanations from both models.
Appendix G Comparison to Rationale-Based Models
We compare B-cos LMs to one rationale-based, explain-then-predict BERT model, RGFS-SA (Saha et al., 2023) [^11] https://huggingface.co/Hate-speech-CNERG/Rationale_predictor on HateXplain. This model leverages human rationales as additional supervision during training. As shown in Table 11, although the RGFS-SA model brings improvement over the conventional BERT model, it generates considerably less faithful rationales compared to B-cos explanations.
Model Method Accuracy ($\uparrow$) Comp ($\uparrow$) Suff ($\downarrow$) SeqPG ($\uparrow$)
Conv. BERT Attention 80.77 22.64 13.83 50
RGFS-SA BERT Rationale 80.09 36.11 16.54 50
B-cos BERT B-cos 78.64 59.66 -4.89 77.57
Table 11: Performance of conventional, RGFS-SA and B-cos (B=1.5) BERT models on HateXplain. SeqPG is consistently 50 for rationale-based models, as their explanations are class-agnostic, similar to attention. The rationale-based RGFS-SA model generates less faithful explanations than B-cos BERT.
Appendix H Human Evaluation Details
@@
In the human study, we select only examples shorter than 25 tokens for HateXplain and 40 tokens for AG News to improve visualization. Additionally, we replace [CLS] and [SEP] with ## to make the examples more understandable for lay users. Below, we provide the instructions along with a detailed description of the criteria and scoring used in our human evaluation. In our human study, 92% of AG News examples and 80% of HateXplain examples contain correct model predictions; in the remaining cases, explanations are supposed to support the wrong predictions.
@@
WARNING: SOME CONTENT IN THIS QUESTIONNAIRE IS HIGHLY OFFENSIVE. Prerequisites: Proficiency in English is required for this evaluation task. If you do not meet this criterion, please do not proceed. We invite you to review 100 examples where LMs perform classification tasks and provide explanations for their predictions. The first 50 examples come from a hate speech detection task, where the model predicts whether a text is toxic or not toxic. The last 50 examples come from a topic classification task, where the model categorizes a text into one of four topics: sports, world, business, or sci/tech. For each example: The model’s prediction is shown along with four explanations justifying the prediction. The order of the explanations is randomized to prevent bias. Words highlighted in green indicate words that had a positive influence on the prediction, while words in red indicate words that had a negative influence. The intensity of the color reflects the strength of the impact. Important: The model’s prediction may be incorrect. Your task is to evaluate the explanations based on how well they support the model’s prediction, not the true labels. Evaluation Task: After reviewing each example, please rate the the human interpretability and human agreement of the four explanations on a scale of 1 to 5. Refer to the definitions and rating scales provided below when making your assessments.
@@
Appendix I More Examples of B-cos Explanations
We provide two more examples of B-cos and other (post-hoc) explanations from AG News in Figure 12. Consistent with our findings in § 4.4, B-cos LMs provide more human interpretable explanations.
@@ Appendix I More Examples of B-cos Explanations
Appendix J Impact of B on Input-weight Alignment
@@
Appendix K Effects of B on Other Metrics
@@
Appendix L B-cos Explanations with Different B Values
@@
Appendix M Example of Model Bias
@@
Appendix N Decoder-Only Model Results
Table 13 presents the average probability gaps between target and foil predictions on every dataset from different vanilla and B-cos models. Table 14 and Table 15 contain MRR results on every dataset in BLiMP and IOI for GPT-2 and Llama-3.2 models, respectively.
Dataset Probability Gap ($\uparrow$)
Vanilla GPT-2 B-cos GPT-2 Vanilla Llama-3.2 B-cos Llama-3.2
aga 0.0120 0.0170 0.0171 0.0196
ana 0.0152 0.0189 0.0140 0.0170
asp 0.0007 0.0008 0.0016 0.0009
dna 0.0011 0.0011 0.0011 0.0017
dnai 0.0021 0.0006 0.0011 0.0013
dnaa 0.0014 0.0012 0.0012 0.0017
dnaai 0.0091 0.0058 0.0077 0.0072
npi 0.0015 0.0002 0.0002 0.0001
darn 0.0067 0.0078 0.0080 0.0085
IOI 0.3351 0.3265 0.4652 0.5021
Table 13: Probability gaps between target and foil next token predictions from vanilla models and B-cos LMs on every dataset.
Dataset Random Grad Norm IxG Occlusion Logit ALTI Logit B-cos
aga 0.6875 0.7927 0.7910 0.7513 0.827 0.964 0.8764
ana 0.7056 0.6753 0.7387 0.5957 0.817 0.976 0.7532
asp 0.3818 0.7512 0.4086 0.4374 0.386 0.499 0.4939
dna 0.4608 0.3629 0.3869 0.9030 0.737 0.646 0.9308
dnai 0.4626 0.4077 0.4317 0.8395 0.711 0.637 0.8596
dnaa 0.4103 0.2632 0.3214 0.6557 0.951 0.807 0.7798
dnaai 0.4074 0.2632 0.3392 0.6167 0.9 0.757 0.7601
npi 0.6121 0.7854 0.4948 0.4775 0.445 0.417 0.4573
darn 0.4888 0.6170 0.3627 0.4247 0.802 0.949 0.8936
IOI 0.2360 0.8599 0.1112 0.8517 1.0 1.0 1.0
Table 14: MRR Alignment of different explanation methods on GPT-2 small predictions on every dataset. B-cos explanations are extracted from the B-cos GPT-2 model. Logit and ALTI Logit results are duplicated from Ferrando et al. (2023).
Dataset Random Grad Norm IxG Occlusion B-cos
aga 0.6868 0.6030 0.5928 0.7811 0.8485
ana 0.7037 0.5955 0.6432 0.6072 0.8535
asp 0.3842 0.7694 0.4537 0.3670 0.6108
dna 0.4615 0.4598 0.4898 0.8352 0.6743
dnai 0.4630 0.4542 0.5043 0.7671 0.6652
dnaa 0.4112 0.4299 0.4750 0.6115 0.6308
dnaai 0.4075 0.4221 0.4498 0.5758 0.5563
npi 0.6123 0.6367 0.7062 0.5154 0.6264
darn 0.4884 0.5828 0.45787 0.5210 0.8065
IOI 0.2328 0.3637 0.1034 0.4767 0.9913
Table 15: MRR Alignment of different explanation methods on Llama-3.2 predictions on every dataset. B-cos explanations are extracted from the B-cos Llama-3.2 model. As Llama models are not supported in Ferrando et al. (2023), we do not include their results.